OpenStudy (angelalee):

Find a polynomial of degree 3 that has zeros 1, -2, and 3, and in which the coefficient of x^2 is 3.

OpenStudy (anonymous):

(x-1)(x+2)(x-3) = 0

OpenStudy (anonymous):

expand and multiply the result by the number that gives a coefficient of '3' on the x^2 term...

OpenStudy (angelalee):

How do you get the coefficient of x^2 to be 3 though?

OpenStudy (anonymous):

x^3 -2x^2 -5x +6 = 0 (-3/2)(x^3 -2x^2 -5x +6 = 0)

OpenStudy (angelalee):

Where did you get the -3/2 from?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

-2*-3/2 =3

OpenStudy (angelalee):

I got up to that part and then i got stuck

OpenStudy (angelalee):

Oh okay thank you so much!

OpenStudy (anonymous):

sure:)